Consider a step-index fiber for which n1 = 1.475, n2 = 1.460 and a = 25µm. What is the maximum value of θ for which the ray ill be guided by the fiber. 

I need this answer sir



1
Expert's answer
2021-05-04T18:47:13-0400

Given,

"n_1=1.475\\n_2=1.460"

"a=25 \ \mu m"

Numerical aperture (NA) ="\sqrt{n_1^2-n_2^2}=\sqrt{(1.475)^2-(1.460)^2}=\sqrt{2.25-2.19}=0.210"

"\theta_{max}=sin^{-1}(0.210)=12.12\degree"